Q. 9 The following is a partial regression result of a two-variable model (i.e. simple linear...

Q. 9

The following is a partial regression result of a two-variable model (i.e. simple linear regression). In the study, a health care economist seeks to determine if a relationship exists between personal income and expenditures on health care, both measured in billions of dollars.

Regression Statistics
Multiple R ???
R Square ???
Standard Error
Observations 51
ANOVA
df SS MS F P-value
Regression 1 15,750.32 0.00001
Residual/Error
Total ??? 16,068.21
Coefficients Standard Error t Stat P-value Lower 95% Upper 95%
Intercept 0.1764964 0.467509347 0.377524859 0.70741 -0.76 1.116
Income 0.1416522 0.002875 ??? 0.00001 ??? 0.147

The sum of squares total (SST) is a measure of ___ and is equal to:

a. Explained variation of Y; 16,068.2143

b. Total variation of Y; 15,750.3157

c. Unexplained variation; 317.8986

d. Total variation of Y; 16,068.2143

e. None of the above

The estimated variance of the residual (error term) is:

a. 55.1830

b. 2.5471

c. 6.4877

d. none of the above

The correlation coefficient of the above model is:

a. 0.99

b. 0.9802

c. 0.9718

d. none of the above
Answer #1

1)

d. Total variation of Y; 16,068.2143   {see the column SS and row Total}

2)

This is given by MSE

= SSE/(n-2) = (SST - SSR)/(n-2)

= ( 16,068.2143 - 15,750.32 )/49

= 6.48763877551

option C)

3)

r = sqrt(SSR/SST)

= sqrt(15,750.32/ 16,068.2143 )

= 0.99005

A)0.99

Please rate